# Inference for Regression with

### 23 Feb 2024 v1

@@ Appendix B Further Details on the Simulation Exercise
To investigate the impact of $\kappa$ on the estimation of $\gamma$, we ran three sets of simulations which vary only by the total number of features drawn per observations. For simplicity, in each of the set of simulations we set $C_{i}$ to be equal for all $i$. We set $C\in\{20,80,160\}$ which, given that $N$ is fixed to 6400 corresponds to $\kappa\in\{4,1,0.5\}$.
@@
We simulated data 40 times for each set and then estimated the model using 1-step approach, 2-step approach and the infeasible 2-step approach with known $\theta$.
@@
Parameter Value Description (a) Data Simulation $N$ $6400$ Number of observations $V$ $300$ Number of distinct features $C_{i}$$\{20,80,160\}$ Total number of features per document $K$ $2$ Number of latent types True $\phi$ 1 Effect of a covariates on un-normalized type shares True $\gamma$ 5 Effect of topic shares on numerical outcomes True $\alpha$ $(0,1,1,1)$ Effect of additional covariates on numerical outcomes $g_{i}$ $\sim N(0,\frac{\log(3)}{1.96})$ Covariate affecting type shares $q_{i,m}\enspace\forall{m\in(1,2,3)}$ $\sim N(0,3)$ Additional covariates affecting outcome $\sigma^{2}_{Y}$ $16$ SD of the numeric outcome’s residual $\sigma^{2}_{\theta}$ $1$ SD of residual of the un-normalized type shares $\eta$ $0.2$ Dirichlet concentration parameter (a) Hyperparameters $K$ as above Number of latent types $\eta$ as above Dirichlet concentration parameter $\sigma^{2}_{\theta}$ as above SD of residual of the un-normalized type shares $p(\phi_{1})$ $N(0,4)$ Prior for $\phi_{1}$, i.e. $\sigma^{2}_{\phi}=4$ $p(\gamma_{1})$ $N(0,100)$ Prior for $\gamma_{1}$, i.e. $\sigma^{2}_{\gamma}=100$ $p(\alpha)\enspace\forall m\in(0,1,2,3)$ $N(0,100)$ Prior for $\alpha$, i.e. $\sigma^{2}_{\alpha}=100$ $p(\sigma_{Y})$ $\operatorname{Gamma}(1,10)$ Prior for $\sigma_{Y}$, i.e. $s_{0}=1$ and $s_{1}=10$
We performed the simulation on a ‘N1-highmem-2’ instance on the Google Could Platform. The instance has 2 vCPUs and 13 GB of memory. We also utilized a single Tesla V100 GPU. We run chose 2000 warmup ans 2000 post-warmup iterations. A single simulation (consisting of drawing the data, and estimating the model in three ways) took approximately 10 minutes.

### 24 Mar 2024 v2

@@ Appendix B Further Details on the Simulation Exercise
To investigate the impact of $\kappa$ on the estimation of $\gamma$, we ran three sets of simulations which vary only by the total number of features drawn per observations. For simplicity, in each of the set of simulations we set $C_{i}$ to be equal for all $i$. We set $C\in\{10,25,200\}$ which, given that $N$ is fixed to $10000$ corresponds to $\kappa\in\{10,4,0.5\}$.
@@
We simulated data 200 times for each set and then estimated the model using 1-step approach, 2-step approach and the infeasible 2-step approach with known $\theta$.
@@
Parameter Value Description (a) Data Simulation $N$ $10000$ Number of observations $V$ $300$ Number of distinct features $C_{i}$$\{10,25,200\}$ Total number of features per document $K$ $2$ Number of latent types True $\phi$ 1 Effect of a covariates on un-normalized type shares True $\gamma$ 5 Effect of topic shares on numerical outcomes True $\alpha$ $(0,1,1,1)$ Effect of additional covariates on numerical outcomes $g_{i}$ $\sim N(0,\frac{\log(3)}{1.96})$ Covariate affecting type shares $q_{i,m}\enspace\forall{m\in(1,2,3)}$ $\sim N(0,3)$ Additional covariates affecting outcome $\sigma^{2}_{Y}$ $16$ SD of the numeric outcome’s residual $\sigma^{2}_{\theta}$ $1$ SD of residual of the un-normalized type shares $\eta$ $0.2$ Dirichlet concentration parameter (b) Hyperparameters $K$ as above Number of latent types $\eta$ as above Dirichlet concentration parameter $\sigma^{2}_{\theta}$ as above SD of residual of the un-normalized type shares $p(\phi_{1})$ $N(0,4)$ Prior for $\phi_{1}$, i.e. $\sigma^{2}_{\phi}=4$ $p(\gamma_{1})$ $N(0,100)$ Prior for $\gamma_{1}$, i.e. $\sigma^{2}_{\gamma}=100$ $p(\alpha)\enspace\forall m\in(0,1,2,3)$ $N(0,100)$ Prior for $\alpha$, i.e. $\sigma^{2}_{\alpha}=100$ $p(\sigma_{Y})$ $\operatorname{Gamma}(1,10)$ Prior for $\sigma_{Y}$, i.e. $s_{0}=1$ and $s_{1}=10$
We performed the simulation on a ‘N1-highmem-2’ instance on the Google Could Platform. The instance has 2 vCPUs and 13 GB of memory. We also utilized a single Tesla V100 GPU. We run chose 500 warmup and 500 post-warmup iterations. A single simulation (consisting of drawing the data, and estimating the model in three ways) took approximately 6 minutes.
B.1 Additional Results
Figure B.1 below, replicates Figure 1 but adds the bias-corrected estimates for $\gamma_{1}$. The bias-corrected estimates are obtained by subtracting the estimated bias computed using the formula from Theorem 1 from the two-step estimates. Since the estimates of the asymptotic bias are negative, the bias-corrected estimates are larger than the two-step estimates. Since the theorem relies on a first-order approximation, the bias-correction works best for small values of $\kappa$. As $\kappa$ increases, the accuracy of the bias-corrected estimates diminishes, and for $\kappa=10$, they perform worse than the uncorrected two-step estimates.
Figure B.1: Evolution of Bias in Regression Coefficients across $\kappa$ Values, including Bias-Corrected Estimates
Appendix C Example Code
Figure C.1: Numpyro’s code used to estimate Supervised Topic Model with Covariates
[…]
import numpyro.distributions as dist
import jax.numpy as jnp
from jax.nn import softmax
[…]
def __init__(self, K, N, V, z, q, eta = .1, alpha = 1):
self.K = K # number of latent types
self.N = N # number of observations
self.V = V # number of distinct features
self.z = z # number of covariates affecting outcome
self.q = q # number of covariates affecting type shares
self.eta = eta
self.alpha = alpha
[…]
# Supervised topic model with covariates
[…]
# X : feature count matrix
# C : total number of features per observation
# Z : covariates entering regression
# Q : covariates entering type shares
# K : number of types
# eta, alpha : Dirichlet hyperparamters
[…]
beta = sample(”beta”, dist.Dirichlet(
self.eta * jnp.ones(self.V - self.num_anchors_per_class)))
[…]
A = sample(”A”, dist.Normal(jnp.matmul(Q, phis) , self.alpha))
[…]
theta = deterministic(
%**** 0file.tex Line 1775 ****”theta”,
softmax(jnp.hstack([A, jnp.zeros([self.D, 1])]), axis = -1)
[…]
total_count=C,
probs = jnp.matmul(theta, beta)
with plate(”hist”, self.N):
X_bows = sample(”obs_x”, distMultinomial, obs = X)
[…]
zetas = sample(”zetas”, dist.Normal(0, 10).expand([self.z]))
sigma = sample(”sigma”, dist.Gamma(1, 10))
[…]
Y = sample(”obs_y”, dist.Normal(mean, sigma), obs = Y)